# Unsupervised Prior Learning: Discovering Categorical Pose Priors from Videos

### v1

@@ 5 Discussion
We introduce the challenge of unsupervised prior learning and highlight its significance in pose estimation. To address this, we propose a novel method called Pose Prior Learner (PPL). PPL utilizes a hierarchical memory to store compositional parts of learnable prototypical poses, which are distilled into a general pose prior for any object category. Our experimental results show that PPL requires no additional knowledge and outperforms recent competitive baselines in video pose estimation. Notably, the learned prior proves to be even more effective in pose estimation than methods that rely on human-defined priors. With hierarchical memory and learned priors, PPL can perform iterative inferences and robustly estimate poses in occluded scenes.
Despite outstanding performance in unsupervised pose estimation, PPL has several limitations. For instance, it requires video data captured in static backgrounds. Additionally, PPL learns 2D priors, which makes it difficult to capture real-world 3D postures. Thus, PPL struggles in scenarios where foreground objects involve rotations or significant shape changes. Extending PPL to incorporate 3D priors will be a key focus of our future research.
@@
We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 64, training for 50 epochs. Unless specified, all images are resized to $256\times 256$. The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero. We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in (He et al., 2022a). For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ tokens of dimension $512$, for all experiments.
@@
Taichi (Siarohin et al., 2019) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances. Following the approach in (Siarohin et al., 2021), we use 5,000 frames for training and 300 frames for testing.
YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL in non-human species. We used only 20 videos from YouTube and decomposed them into a total of 2000 video frames for training. We use this customized dataset to demonstrate the applicability of PPL in learning pose priors for non-human categories. All images are trained and tested at a resolution of $256\times 256$. We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos.
Appendix A3 Ablation on Memory Bank Tokens and Number of Keypoints
Figure A1: Ablation of our PPL method on Memory Bank Tokens and Number of Keypoints in Human3.6m. The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank tokens (ranging from 64 to 512). The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank tokens.

### v2

@@ 5 Discussion
We introduce the challenge of unsupervised categorical prior learning and highlight its significance in pose estimation. To address this, we propose a novel method called Pose Prior Learner (PPL). PPL utilizes a hierarchical memory to store compositional parts of learnable prototypical poses, which are distilled into a general pose prior for any object category. Our experimental results show that PPL requires no additional human annotations and outperforms recent competitive baselines in pose estimation. Notably, the learned prior proves to be even more effective in pose estimation than methods that rely on human-defined priors. With hierarchical memory and learned priors, PPL can perform iterative inferences and robustly estimate poses in occluded scenes. Despite outstanding performance in unsupervised categorical pose estimation, PPL has several limitations. For instance, it learns 2D priors, which makes it difficult to capture real-world 3D postures. Thus, PPL struggles in scenarios where objects involve rotations or significant shape changes. Extending PPL to incorporate 3D priors will be a key focus of our future research.
6 Impact Statement
We present the challenge of unsupervised learning of categorical priors for pose estimation and propose Pose Prior Learner (PPL) to address it. By leveraging learned priors, PPL enhances robustness and generalizability during inference, promoting the safe deployment of AI technologies. However, privacy concerns surrounding images and videos in these datasets, especially for human pose estimation, must be carefully addressed. Protecting personal and sensitive information remains paramount as AI advances.
@@ 6 Impact Statement
We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 64, training for 50 epochs. Unless specified, all images are resized to $256\times 256$. The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero. We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in (He et al., 2022a). For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$, for all experiments.
@@
Taichi (Siarohin et al., 2019b, a) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances. Following the approach in (Siarohin et al., 2021), we use 5,000 frames for training and 300 frames for testing.
YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL. Existing dog datasets are not suitable, as they often contain multiple, partially occluded dog instances. Therefore, we curated a custom dataset using 20 YouTube videos, extracting 2,000 frames for training. This dataset allows us to demonstrate PPL’s ability to learn pose priors for non-human categories without using ground truth poses. All images are trained and tested at a resolution of $256\times 256$. We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos. The dataset will be publicly released together with other data, models, and source code.
CUB-200-2011 (Wah et al., 2011) contains 11788 images of birds. We crop and align the images according to (He et al., 2022a). We use the train/val/test split of (Choudhury et al., 2021). All images are trained and tested at a resolution of $128\times 128$.
11k-Hands (Afifi, 2019), Horse (Zhu et al., 2017), and Flowers (Nilsback & Zisserman, 2008) are used for qualitative visualization. Images with multiple horses are removed and all horses are aligned to face left. All images are trained and tested at a resolution of $128\times 128$.
Appendix A3 Ablation on Memory Bank Vectors and Keypoint Numbers
Figure A0: Ablation of our PPL method on memory bank vectors and number of keypoints in Human3.6m. The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank vectors (ranging from 64 to 512). The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank vectors.
Appendix A4 PPL Results of Keypoint Detection on Occluded Images
Figure A1: PPL results of keypoint detection as a function of number of inference iterations on images with RandomMasking from Human3.6m. The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
Figure A2: PPL results of keypoint detection as a function of number of inference iterations on images with CenterMasking from Human3.6m. The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.
Appendix A5 Retrieval and Distillation of the proposed Hierarchical Memory in PPL
Appendix A6 Training Techniques
To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017). Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step. (2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy. This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time. (3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$. This creates a chicken-and-egg problem that complicates training. To address this, we introduce two gradient detachments to separate the training processes. First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2. Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure ALABEL:fig:mem_stru in Appendix A5.
Appendix A7 Additional Visualizations